# Digital requirements engineering with an INCOSE-derived SysML meta-model

### v1

@@ Abstract
Traditional requirements engineering tools do not readily access the SysML-defined system architecture model, often resulting in ad-hoc duplication of model elements that lacks the connectivity and expressive detail possible in a SysML-defined model. Further integration of requirements engineering activities with MBSE contributes to the Authoritative Source of Truth while facilitating deep access to system architecture model elements for V&V activities. We explore the application of MBSE to requirements engineering by extending the Model-Based Structured Requirement SysML Profile to comply with the INCOSE Guide to Writing Requirements while conforming to the ISO/IEC/IEEE 29148 standard requirement statement patterns. Rules, Characteristics, and Attributes were defined in SysML according to the Guide to facilitate requirements definition, verification & validation. The resulting SysML Profile was applied in two system architecture models at NASA Jet Propulsion Laboratory, allowing us to assess its applicability and value in real-world project environments. Initial results indicate that INCOSE-derived Model-Based Structured Requirements may rapidly improve requirement expression quality while complementing the NASA Systems Engineering Handbook checklist and guidance, but typical requirement management activities still have challenges related to automation and support in the system architecture modeling software.
Keywords digital engineering (DE); model-based systems engineering (MBSE); model-based structured requirement (MBSR); requirements engineering; verification & validation (V&V); INCOSE; NASA
@@ 1 Introduction
As engineered systems grow in complexity, the demand for more cost-effective system development programs grows in turn. A critical point of leverage in reducing system development costs is by improving Requirements Engineering (RE) processes and the quality of its outputs (Hirshorn et al., 2017). Digital Engineering (DE) promises to improve quality and reduce costs through increased access and connectivity of digital artifacts in a central data store called the Authoritative Source of Truth (ASoT) (Noguchi et al., 2020). Model-Based Systems Engineering (MBSE) improves upon Document-Centric Systems Engineering (DCSE) by incorporating formal digital models throughout systems engineering processes and products, by leveraging those digital system models to precisely represent design values and relationships, and by computing model validity based on modeling language and design rules. MBSE practice develops Digital RE (DRE) as a response to increasing the system architecture model connectivity toward complete traceability of stakeholder needs and system requirements in support of DE goals.
The effectiveness of MBSE compared to DCSE is now well-studied in the literature (Rogers III and Mitchell, 2021; Younse et al., 2021; Madni and Purohit, 2019; Carroll and Malins, 2016) making it the most important systems engineering practice toward achieving DE goals in the organization. As a motivating example, the NASA-ESA Mars Sample Return mission is “an ambitious and complex space system engineering endeavor” (Sundararajan, 2022) with multiple interfacing space systems necessary to coordinate the safe return of Martian gas and solid core samples for further study on Earth. The MSR mission has recently undergone its second Independent Review Board assessment that includes a probable program life cycle cost estimate of US$8-11 billion, “strong irrefutable evidence” that strong systems engineering (SE) is a crucial factor for mission success, and recommendations to refactor the program architecture to control costs (Figueroa et al., 2023). Strong SE is now model-based, and DRE supports the effectiveness of SE by improving the quality of requirements early on through the application of model-based rules and through precise definition of model traceability; and it helps reduce costs through model-based generation of stakeholder-tailored views and through reduced rework due to inadequate requirements expression. DRE is not a substitute for good RE practice, but rather it is a vision for a model-based and digitally-augmented RE discipline still maturing from its DCSE roots.
@@ 1.1 Digital Engineering
Recognition of the current and potential impact of digital models, including those used in MBSE, has led to the development by the US Department of Defense of a strategy for taking greater advantage of digital models to transform the system development process. DE is “an integrated digital approach that uses authoritative sources of system data and models as a continuum across disciplines to support lifecycle activities from concept through disposal” (Office of the Deputy Assistant Secretary of Defense for Systems Engineering, 2018). As an engineering leader, NASA has invited the future of digital workflows by publishing a Digital Transformation strategy (Marlowe et al., 2022), MBSE strategy (Weiland, 2021), and DE Acquisition Framework Handbook (Office of the NASA Chief Engineer, 2020). DE is not a new discipline of engineering but rather an intentional transformation of how an organization integrates and performs its engineering activities to achieve higher quality and efficiency (Noguchi et al., 2020).
One of the DE goals is to provide an enduring ASoT of the system to improve communication and decision-making. The system architecture model is one component of the ASoT, typically integrated in a centralized repository, and the system requirements may be created in the architecture model or synchronized with the model from a Requirement Management Tool (RMT). DRE further integrates requirements with the ASoT, enabling formal verification and validation (V&V) activities that may be automated to improve model confidence and ease stakeholder reviews (Duprez et al., 2023).
@@ 1.2 Requirements Engineering
RE is a subset of systems engineering that encompasses requirements development and requirements management. The ISO/IEC/IEEE 29148:2018 standard defines requirements engineering as “an interdisciplinary function that mediates between the domains of the acquirer and supplier or developer to establish and maintain the requirements to be met by the system, software or service of interest. Requirements engineering is concerned with discovering, eliciting, developing, analyzing, verifying (including verification methods and strategy), validating, communicating, documenting and managing requirements” (ISO/IEC/IEEE, 2018). The range of RE activities necessitates the use of metadata to organize information about each requirement, emphasizing that the familiar “shall” statement is only one attribute of a well-managed and model-connected requirement.
The INCOSE Guide to Writing Requirements (GtWR) provides a current perspective of well-formed requirements, and it defines a requirement statement as “the result of a formal transformation of one or more sources, needs, or higher-level requirements into an agreed-to obligation for an entity to perform some function or possess some quality within specified constraints with acceptable risk” (Wheatcraft and Ryan, 2023). The GtWR emphasizes that the requirement statement forms the basis of contractual language, and then presents a rules-based structured format for facilitating that communication. It defines a requirement expression as the requirement statement and its attributes. The GtWR recommends a data-centric practice using a RMT, as opposed to spreadsheets or documents, to model and present requirement expressions using diagrams and tables for stakeholder-tailored views.
Systems engineering handbooks provide another important reference for requirements engineering activities, complementing the detailed guides, manuals, and standards cited above. The INCOSE Systems Engineering Handbook - Fifth Edition provides updated Sections 2.3.5.2 and 2.3.5.3 that incorporate the latest INCOSE guides and manuals on needs and requirements engineering (INCOSE, 2023). The NASA Systems Engineering Handbook (Hirshorn et al., 2017) describes the traditional NASA requirements definition and management processes in Sections 4.2 and 6.2, respectively, emphasizing bidirectional traceability, and including a checklist in Appendix C and an informal set of characteristics similar to those defined in the latest INCOSE GtWR. The INCOSE GtWR and NASA sets of characteristics are compared in Sec. 3 and found to be complementary.
@@ 1.3 Systems Modeling Language
The Systems Modeling Language (SysML) by the Object Management Group (OMG) (OMG, 2022) is a standard language for modeling system architectures which provides the capability to model the solution space as structure, behavior, rules and requirements as digital model elements in a directional graph constrained by SysML semantics. SysML version 1.7 is expected to be the final version in the 1.x series as the standards development effort shifts to the new version 2. SysML provides rudimentary facilities for modeling system requirements, including the primary attributes: ID, name, and text; and the relationships: derive, refine, satisfy, verify, and trace (which is discouraged in favor of the more precise relationships). Requirement type and rationale are not attributes provided by the SysML standard but are customizations of the SysML Profile often provided by systems architecture modeling tools. According to the standard, SysML Requirements may be shown in a Requirements Diagram, or placed on other SysML diagrams to highlight relationships for certain stakeholder views; requirements tables and matrices are non-normative.
@@
Today RE is often practiced with the help of an RMT which stores the requirements in a database and provides structured access to them for management activities. Since this RE practice uses digital models in a computer database, it may be considered MBSE or DE. However, the threat remains of duplicative system model elements interfering with traceability due to the RMT lacking the SysML meta-model used to define the system architecture. Controlled import/export or RMT data connector synchronization cycles alleviate this problem but may cause issues due to the non-standard interfaces between tools (Wheaton and Herber, 2024). Although SysML v2 has addressed this issue by standardizing the API providing access to the system architecture model, it will be years before it sees widespread adoption inside DE organizations (Bajaj et al., 2022). SysML v2 improves upon v1 by modeling requirements as constraints that must be satisfied by the system-of-interest, and it uses the consistent distinction between a requirement definition and requirement usage to aid in reuse (OMG, 2024).
@@
Due to the perceived inadequate facilities for modeling and managing requirements, SysML has not been favored by requirements engineers (Wheatcraft and Ryan, 2023). Its apparent advantages with respect to its graphical syntax (diagrams) have not been enough to satisfy the critical need of managing possibly thousands of requirements. However, SysML makes it possible to extend the language through Stereotypes, thus availing the systems engineer with meta-modeling capabilities to define custom elements and to use them as just another SysML element. This meta-modeling capability is what enables the Model-Based Structured Requirement (MBSR) approach described in this paper.
@@ 1.4 Overview
This paper presents an extension to the MBSR approach developed by Herber and Eftekhari-Shahroudi (2023) and Herber et al. (2022) that incorporates the INCOSE GtWR ontology and the ISO/IEC/IEEE (2018) standard pattern for requirement statements (Fig. 1) using a notional space flight system to demonstrate the use of the SysML Profile. The rest of the paper is organized as follows: Sec. 2 describes how the INCOSE-derived MBSR meta-model is defined and used; Sec. 3 discusses the benefits and challenges of this approach; Sec. 4 compares MBSR to other model-based requirements methods; Sec. 5 discusses some limitations of this current research and identifies areas for future work; and Sec. 6 concludes and reflects on the utility of MBSR in the context of DRE.
2 INCOSE-derived Model-Based Structured Requirement
A strong SE practice is a countervailing force against cost overruns and system development project cancellations, and it is bolstered by a strong embedded RE practice from the beginning of the project. Where there were undefined, unsupported, or undisciplined RE processes in an organization, tailored guidance from INCOSE should fill the knowledge gap. Developing successful complex systems requires participation from diverse sets of stakeholders and organizations, leading to the development of Simplified Technical English (AeroSpace and Defence Industries Association of Europe, 2021) to reduce language ambiguity and confusion. Unlike the situation in the late 20th century when software development projects turned into production messes and accumulated technical debt, practitioners today may now benefit from advancements in RE to minimize ambiguity, reduce technical debt, formalize V&V processes, and produce consistently high-quality design output specifications (Wheatcraft et al., 2022; Avdeenko and Pustovalova, 2016).
This section of the paper describes classical standard SysML requirements modeling and performs a gap analysis (Sec. 2.1), structured requirements and its related work (Sec. 2.2), and INCOSE GtWR-derived meta-model extensions to MBSR from Herber and Eftekhari-Shahroudi (2023) (Sec. 2.3).
@@ 2.1 Classical SysML Requirements Modeling
Classical SysML represents a requirement as an “indivisible entity” (Dick et al., 2017) that enriches and is enriched by other system architecture model elements through the use of typed relationships. According to the SysML standard specification, “a requirement is defined as a stereotype of UML Class subject to a set of constraints” and includes “properties to specify its unique identifier and text requirement,” noting that “additional properties … can be specified by the user” (OMG, 2022). The ‘name’, ‘id’, and ‘text’ properties of SysML requirements are defined as a simple string of characters, lacking any other structure, manipulable only by standard string processing functions of popular programming languages if available. Traceability relationships include containment, and subtypes of the UML Dependency relationship, as follows:
@@
specifies a directed relationship used “to describe how a model element or set of elements can be used to further refine a requirement” (OMG, 2022).
@@
specifies a type of Trace that “describes how a design or implementation model satisfies one or more requirements” (OMG, 2022).
@@
is a general-purpose relationship between a requirement and any other model element; its use with more specific traceability relationships listed above is discouraged due its ambiguity (OMG, 2022).
Classical SysML-based requirements popularized the idea of “cohabitation” where requirements and system model elements exist and are linked together in the same model (Bernard, 2012) (Fig. 2). Practicing DRE with this MBSE approach leads to the notion that requirements engineers and system architects should work in integrated teams rather than organizational silos and separate software tools (Wheatcraft et al., 2022). However, due to the lack of requirements management facilities such as software-defined workflows, audit logs, and change management boilerplate, managing requirements in SysML is considered inferior compared to a purpose-built RMT (Wheatcraft and Ryan, 2023).
@@
DRE has outgrown classical SysML requirements due to their lack of precision, leading practitioners and tool vendors to extend standard SysML in various ways to satisfy MBSE objectives (OMG, 2022). The Trace and Copy relationships are discouraged, leading to confusion about their use, although this may be mitigated by hiding them in the SysML tool’s user perspective. The Verify relationship perpetuates the conflation of system V&V with requirements V&V (“verifying a requirement”), a distinction made clear in the INCOSE GtWR. The non-normative extensions, such as the additional requirement stereotypes and risk kinds in OMG (2022), may confuse practitioners due to the ‘one size fits all’ approach, and interface requirements, in particular, are rejected by the GtWR. Although SysML provides the basic utilities for complex system requirements analysis, the implementation of such analyses can take considerable effort, favoring organizations with mature model libraries. In summary, classical SysML requirements lack modeling precision without significant extensions, and introduce ambiguity and confusion in modeling constructs that contradict the latest RE guidance from INCOSE and ISO/IEC/IEEE.

### v2

@@ Abstract
Traditional requirements engineering tools do not readily access the SysML-defined system architecture model, often resulting in ad-hoc duplication of model elements that lacks the connectivity and expressive detail possible in a Systems Modeling Language (SysML)-defined model. Further integration of requirements engineering activities with MBSE contributes to the Authoritative Source of Truth while facilitating deep access to system architecture model elements for V&V activities. Building from prior work on Model-Based Structured Requirements, we extend its SysML profile to comply with the INCOSE Guide to Writing Requirements while conforming to the ISO/IEC/IEEE 29148 standard requirement statement patterns. Rules, Characteristics, and Attributes were defined in SysML according to the Guide to facilitate requirements development, verification & validation. The resulting SysML profile was applied in two system architecture models at NASA Jet Propulsion Laboratory, allowing us to assess its applicability, value, and gaps in real-world project environments. Initial results indicate that use of the INCOSE-derived Model-Based Structured Requirement profile may rapidly improve requirement expression quality while complementing the NASA Systems Engineering Handbook checklist and guidance, but typical requirement management activities still have challenges related to automation and support in the SysML modeling software.
Keywords digital engineering (DE); model-based systems engineering (MBSE); Systems Modeling Language (SysML); requirements engineering; INCOSE Guide to Writing Requirements (GtWR); NASA/ESA Mars Sample Return (MSR)
@@ 1 Introduction
As engineered systems grow in complexity, the demand for more cost-effective system development programs grows in turn. A critical point of leverage in reducing system development costs is by improving requirements engineering processes and the quality of its outputs (NASA_SE_HB2017). Digital engineering (DE) promises to improve quality and reduce costs through increased access and connectivity of digital artifacts in a central data store called the authoritative source of truth (ASoT) (noguchi2020digital). Model-based systems engineering (MBSE) improves upon document-centric systems engineering by incorporating formal digital models throughout systems engineering processes and products, by leveraging those digital system models to precisely represent design values and relationships, and by computing model validity based on modeling language and design rules. MBSE practice develops digital requirements engineering (DRE) as a response to increasing the system architecture model connectivity toward complete traceability of stakeholder needs and system requirements in support of DE goals.
With MBSE being central to the INCOSE systems engineering vision dating back to 2007 (INCOSEvision2020) and DE later being added in 2014 (INCOSEvision2025), and carried forward to the 2035 Vision (INCOSEvision2035), systems engineers should apply effort to incorporate model-based, digital methods in their practice. The effectiveness of MBSE compared to document-centric systems engineering continues to be studied, but results already corroborate its expected benefits (rogers2021mbse; younse2021mbse; systems7010012; osti_1561164), making it the most important systems engineering practice toward achieving DE goals in the organization. As a motivating example, the NASA-ESA Mars Sample Return campaign is “an ambitious and complex space system engineering endeavor” (sundararajan2022understanding) with multiple interfacing space systems necessary to coordinate the safe return of Martian gas and solid core samples for further study on Earth. The Mars Sample Return program completed its second Independent Review Board assessment that included a probable program life cycle cost estimate of US$8-11 billion, “strong irrefutable evidence” that strong systems engineering (SE) is a crucial factor for mission success, and recommendations to refactor the program architecture to control costs (MSR_IRB2023). Strong systems engineering is increasingly model-based, and while DRE is not a substitute for good requirements engineering practice, it highlights a paradigm shift that includes other digital engineering trends.
@@ 1.1 Digital Engineering
Recognition of the current and potential impact of digital models, including those used in MBSE, has led the US Department of Defense to develop a strategy that makes digital models essential to the system development process. DE is “an integrated digital approach that uses authoritative sources of system data and models as a continuum across disciplines to support lifecycle activities from concept through disposal” (DoD_DE_Strategy2018). As an engineering leader, NASA has invited the future of digital workflows by publishing a Digital Transformation strategy (marlowe2022), MBSE strategy (weiland2021), and DE Acquisition Framework Handbook (NASA_DE_HDBK). DE is not a new discipline of engineering but rather an intentional transformation of how an organization integrates and performs its engineering activities to achieve higher quality and efficiency (noguchi2020digital).
One of the DE goals is to provide an enduring ASoT of the system to improve communication and decision-making. The system architecture model is one component of the ASoT, typically integrated in a centralized repository, and the system requirements may be created in the architecture model or synchronized with the model from a requirements management tool (RMT). DRE further integrates requirements with the ASoT, enabling formal verification and validation (V&V) activities that may be automated to improve model confidence and ease stakeholder reviews (duprez2023approach).
@@ 1.2 Requirements Engineering
Requirements engineering (RE) is a process area of systems engineering that encompasses requirements development and requirements management. The ISO/IEC/IEEE 29148:2018 standard defines requirements engineering as “an interdisciplinary function that mediates between the domains of the acquirer and supplier or developer to establish and maintain the requirements to be met by the system, software or service of interest. Requirements engineering is concerned with discovering, eliciting, developing, analyzing, verifying (including verification methods and strategy), validating, communicating, documenting and managing requirements” (ISO29148). The range of RE activities necessitates the use of metadata to organize information about each requirement, emphasizing that the familiar “shall” statement is only one attribute of a well-managed and model-connected requirement.
The INCOSE Guide to Writing Requirements (GtWR) provides a current perspective of well-formed requirements, and it defines a requirement statement as “the result of a formal transformation of one or more sources, needs, or higher-level requirements into an agreed-to obligation for an entity to perform some function or possess some quality within specified constraints with acceptable risk” (GtWR2023). The GtWR emphasizes that the requirement statement forms the basis of contractual language, and then presents a rules-based structured format for facilitating that communication. It defines a requirement expression as the requirement statement and its attributes. The GtWR recommends a data-centric practice using a RMT, as opposed to spreadsheets or documents, to model and present requirement expressions using diagrams and tables for stakeholder-tailored views.
Systems engineering handbooks provide another important reference for requirements engineering activities, complementing the detailed guides, manuals, and standards cited above. The INCOSE Systems Engineering Handbook - Fifth Edition provides updated Sections 2.3.5.2 and 2.3.5.3 that incorporate the latest INCOSE guides and manuals on needs and requirements engineering (INCOSE_SE_HB2023). The NASA Systems Engineering Handbook (NASA_SE_HB2017) describes the traditional NASA requirements definition and management processes in Sections 4.2 and 6.2, respectively, emphasizing bidirectional traceability, and including a checklist in Appendix C and an informal set of characteristics similar to those defined in the latest INCOSE GtWR. The INCOSE GtWR and NASA sets of characteristics are compared in Sec. 4 and found to be complementary.
@@ 1.3 Systems Modeling Language
The Systems Modeling Language (SysML) by the Object Management Group (OMG) (SysML1.7) is a standard language for modeling system architectures which provides the capability to model the solution space as structure, behavior, rules and requirements as digital model elements in a directional graph constrained by SysML semantics. SysML version 1.7 is expected to be the final version in the 1.x series as the standards development effort shifts to the new version 2. SysML provides rudimentary facilities for modeling system requirements, including the primary attributes: ID, name, and text; and the relationships: derive, refine, satisfy, verify, and trace (which is discouraged in favor of the more precise relationships). Requirement type and rationale are not attributes provided by the SysML standard but are customizations of the SysML profile often provided by systems architecture modeling tools. According to the standard, SysML Requirements may be shown in a Requirements Diagram, or placed on other SysML diagrams to highlight relationships for certain stakeholder views; requirements tables and matrices are non-normative.
@@
Today RE is often practiced with the help of an RMT which stores the requirements in a database and provides structured access to them for management activities. Since this RE practice uses digital models in a computer database, it may be considered MBSE or DE. However, the threat remains of duplicative system model elements interfering with traceability due to the RMT lacking the SysML meta-model used to define the system architecture. Controlled import/export or RMT data connector synchronization cycles alleviate this problem but may cause issues due to the non-standard interfaces between tools (Wheaton2024a). Although SysML v2 has addressed this issue by standardizing the API providing access to the system architecture model (bajaj2022sysmlv2), due to disparate tool development lifecycles and enterprise adoption timelines, it will be years before it sees widespread adoption inside organizations (Call2022b). SysML v2 improves upon v1 by modeling requirements as constraints that must be satisfied by parts of the system-of-interest, and it uses the consistent distinction between a requirement definition and requirement usage to aid in reuse (SysMLv2).
@@
Due to the perceived inadequate facilities for modeling and managing requirements using SysML, the GtWR cautions against its use (GtWR2023). Its apparent advantages with respect to its graphical syntax (diagrams) and traceability support have not been enough to satisfy the critical needs of managing and distributing possibly thousands of requirements. In addition, due to accessibility and training challenges associated with SysML modeling tools, the GtWR notes that stakeholders “will still prefer and demand” electronic/printed documents “for the foreseeable future”. However, SysML makes it possible to extend the language through Stereotypes, thus availing the systems engineer with meta-modeling capabilities to define custom elements and to use them to model RE concepts directly. SysML tools also provide means to generate artifacts such as requirements documents from the architecture model, ameliorating the need for specialized software for stakeholder review. This SysML meta-modeling capability specifically enables the GtWR-derived Model-Based Structured Requirement (MBSR) approach described in this paper, while model-based artifact generation enabled benefits and encountered challenges discussed in Section 4.
@@ 1.4 Overview
This paper presents an extension to the MBSR approach developed by herber2023 and herber2022model that incorporates the INCOSE GtWR ontology and the ISO29148 standard pattern for requirement statements (Fig. 4), using a notional space flight system to demonstrate the use of the SysML profile. Model-based application of the GtWR represents a gap in the literature that this paper aims to address.
The contributions of this paper may be summarized as follows:
We present a complete open-source SysML profile for digital requirements engineering based on structured requirement statements from ISO/IEC/IEEE 29148:2018 and formalization of the meta-model from INCOSE GtWR, not available from INCOSE nor similarly published by other research we reviewed.
We demonstrate the application of the INCOSE-derived MBSR profile using an example model in the aerospace domain, adding to the domains studied using MBSR, and describing additional supported capabilities with respect to DRE practice.
We validate INCOSE-derived MBSR in a real-world NASA/ESA Mars Sample Return project in Pre-Phase A (NASA_SE_HB2017) and discuss the perceived benefits and challenges. This work shows how MBSR was applied early in the concept system lifecycle phase compared to late in the development phase which was covered in prior MBSR work. Readers in the aerospace domain may examine this evidence when considering to apply INCOSE-derived MBSR in their work.
The rest of the paper is organized as follows: Sec. 2 describes the prior MBSR SysML language extension that connects model elements to their respective statement pattern slots; Sec. 3 extends the prior MBSR with our INCOSE-derived SysML meta-model and presents a representative example and summarized case study; Sec. 4 discusses the benefits and challenges of this approach compared to unstructured document-based and classical SysML requirements in the context of a real project in Pre-Phase A; Sec. 5 presents a literature review and compares SysML- and other model-based requirement techniques to INCOSE-derived MBSR; and Sec. 6 concludes with limitations and future work, and reflects on the value in the context of DRE.
2 Model-Based Structured Requirement (MBSR)
A strong SE practice is a countervailing force against cost overruns and system development project cancellations, and it is bolstered by a strong embedded RE practice from the beginning of the project. Where there were undefined, unsupported, or undisciplined RE processes in an organization, tailored guidance from INCOSE should fill the knowledge gap. Developing successful complex systems requires participation from diverse sets of stakeholders and organizations, leading to the development of Simplified Technical English (ASDSTE100) to reduce language ambiguity and confusion. Unlike the situation in the late 20th century when software development projects turned into production messes and accumulated technical debt, practitioners today may now benefit from advancements in RE to minimize ambiguity, reduce technical debt, formalize V&V processes, and produce consistently high-quality design output specifications (NRM2022; avdeenko2016ontology).
This section of the paper describes classical SysML requirements modeling and performs a gap analysis (Sec. 2.1), structured requirements and its related work (Sec. 2.2), and INCOSE GtWR-derived meta-model extensions to MBSR from herber2023 (Sec. 2).
@@ 2.1 Classical SysML Requirements Modeling
Classical SysML represents a requirement as an “indivisible entity” (dick2017integrating) that enriches and is enriched by other system architecture model elements through the use of typed relationships. According to the SysML standard specification, “a requirement is defined as a stereotype of UML Class subject to a set of constraints” and includes “properties to specify its unique identifier and text requirement,” noting that “additional properties … can be specified by the user” (SysML1.7). The ‘name’, ‘id’, and ‘text’ properties of SysML requirements are defined as a simple string of characters, lacking any other structure, manipulable only by standard string processing functions of popular programming languages if available. Traceability relationships include containment, and subtypes of the UML Dependency relationship, as follows:
@@
specifies a directed relationship used “to describe how a model element or set of elements can be used to further refine a requirement” (SysML1.7).
@@
specifies a type of Trace that “describes how a design or implementation model satisfies one or more requirements” (SysML1.7).
@@
is a general-purpose relationship between a requirement and any other model element; its use with more specific traceability relationships listed above is discouraged due its ambiguity (SysML1.7).
Classical SysML-based requirements popularized the idea of “cohabitation” where requirements and system model elements exist and are linked together in the same model (bernard2012requirements) (Fig. 1). Practicing DRE with this MBSE approach leads to the notion that requirements engineers and system architects should work in integrated teams rather than organizational silos and separate software tools (NRM2022). However, due to the lack of requirements management facilities such as software-defined workflows, audit logs, and change management boilerplate, managing requirements in SysML presents additional challenges already addressed by enterprise-grade RMTs. While the GtWR notes that “SysML requirement diagrams … are not well-suited to representing multiple or large sets of requirements” (GtWR2023), we find that the same may also apply to requirement tables compared to RMTs (see Sec. 4.2).
@@
DRE has outgrown classical SysML requirements due to their lack of precision, leading practitioners and tool vendors to extend standard SysML in various ways to satisfy MBSE objectives (SysML1.7). The Trace and Copy relationships are discouraged, leading to confusion about their use, although this may be mitigated by hiding them in the SysML tool’s user perspective. The Verify relationship perpetuates the conflation of system V&V with requirements V&V (“verifying a requirement”), a distinction made clear in the INCOSE GtWR. The non-normative extensions, such as the additional requirement stereotypes and risk kinds in SysML1.7, may confuse practitioners due to the ‘one size fits all’ approach, and interface requirements, in particular, are rejected by the GtWR. Although SysML provides the basic utilities for complex system requirements analysis, the implementation of such analyses can take considerable effort, favoring organizations with mature model libraries. In summary, classical SysML requirements lack modeling precision without significant extensions, and introduce ambiguity and confusion in modeling constructs that contradict the latest RE guidance from INCOSE and ISO/IEC/IEEE.